OpenStudy (anonymous):

How do I make this number sentence true: -3+5*2-(-6)=16

OpenStudy (anonymous):

if it's not true, you can't make it . !...

OpenStudy (ash2326):

\[-3+5*2-(-6)\ne16\] to make it true add 3 to the right hand side \[-3+10+6+3=16\]

OpenStudy (anonymous):

@ash2326 that's not allowed O_O

OpenStudy (ash2326):

Why@Kreshnik ?

OpenStudy (anonymous):

you can't add only to one side... this is what i've learned at school . :( ...

OpenStudy (ash2326):

He's asking to make it true, We add to both sides when both sides are equal:)

OpenStudy (experimentx):

well if it's a trick ... then anything is allowed.

OpenStudy (anonymous):

..as @experimentX said, if so.. then ok, but why do I see a "=" sing there?

OpenStudy (experimentx):

it's just a trick .... there's no way 2 = 1

OpenStudy (ash2326):

We do see that it's false expression, so gotta add or subtract something to make it true::P

OpenStudy (anonymous):

if this was the question ..how do we make this equal: \[\Large -3+5\cdot 2-(-6)\neq 16\] I would use that too :)

OpenStudy (experimentx):

i agree your logic.

OpenStudy (anonymous):

anyway , what asker is asking for, I think ash2326 is right... so.. ;)

OpenStudy (experimentx):

it must be trick ...

OpenStudy (anonymous):

I figured it out: (-3 +5) * (2 -(-6)) = 16

OpenStudy (ash2326):

Oh we never thought like that:) Great work @rjgonzo

OpenStudy (anonymous):

whoaaa.... :/

OpenStudy (experimentx):

was a bracket trick ,,, to be exact